At a certain temperature, this reaction establishes an equillbrkum with the given equililbrium constant, Ke K, 2.13 x 10 If, at this temperature, 1.20 mol of A and 3.70 mol of B are placed in a 1.00-L container, what are the concentrations of A, B, and C at equilibrium? Number Number [B Number MI
Answer #1

Here, A is the limiting reagent, as it will get consumed completely and B will still be remaining. Therefore, A will be 0 at equilibrium. Now, we should find how much B has reacted.

3A + 2B = 4C

A + 2/3 B = 4/3 C

1.2A + 0.8B = 1.6C

Thus, 0.8 moles of B has reacted, and 3.7-0.8 = 2.9 moles of B remained at equilibrium.

[A] = 0 M

[B] = 2.9 M

[C] = 1.6 M